Please help me solve this problem of System of equation by substitution.
3x+2y=11
2x-y=7
i need to solve for X and y
I also need to solve a problem of system of equation by Elimination.
5x-y=13
2x-3y=12
i need to solve for X and Y

Answer by checkley77(12844)

3x+2y=11
2x-y=7 OR -Y=-2X+7 OR Y=2X-7
NOW REPLACE Y IN THE FIRST EQUATION WITH (2X-7)
3X+2(2X-7)=11
3X+4X-14=11
7X=11+14
7X=25
X=25/7 ANS.
2*25/7-Y=7
50/7-Y=7
-Y=7-50/7
-Y=49/7-50/7
-Y=-1/7
Y=1/7 ANS.